Calculate total installment payments

In 2018, NB Inc.'s federal taxable income was $242,000. Compute the required installment payments of 2019 tax in each of the following cases:
 

  1. NB’s 2019 taxable income is $593,000.

  2. NB’s 2019 taxable income is $950,000.

  3. NB’s 2019 taxable income is $1,400,000.

Calculate the total installment payments
Answer #1
Answer




a) Total installment payment$     2,01,620

b) Total installment payment$     3,23,000

c) Total installment payment$     4,76,000




Explanation :
a)113900+34% (593000-335000)

= 113900+87720

= $201,620



b)113900+34% (950000-335000)

= 113900+209100

= $323,000



c)113900+34% (1400000-335000)

= 113900+362,100

= $476,000



answered by: DayBuggy